# Backdoor Attack on Vertical Federated Graph Neural Network Learning

### v1

@@ Abstract
Federated Graph Neural Network (FedGNN) is a privacy-preserving machine learning technology that combines federated learning (FL) and graph neural networks (GNNs). It offers a privacy-preserving solution for training GNNs using isolated graph data. Vertical Federated Graph Neural Network (VFGNN) is an important branch of FedGNN, where data features and labels are distributed among participants, and each participant has the same sample space. Due to the difficulty of accessing and modifying distributed data and labels, the vulnerability of VFGNN to backdoor attacks remains largely unexplored. In this context, we propose BVG, the first method for backdoor attacks in VFGNN. Without accessing or modifying labels, BVG uses multi-hop triggers and requires only four target class nodes for an effective backdoor attack. Experiments show that BVG achieves high attack success rates (ASR) across three datasets and three different GNN models, with minimal impact on main task accuracy (MTA). We also evaluate several defense methods, further validating the robustness and effectiveness of BVG. This finding also highlights the need for advanced defense mechanisms to counter sophisticated backdoor attacks in practical VFGNN applications.
Keywords FedGNN  $\cdot$ Backdoor Attack  $\cdot$ Graph Neural Network
@@ 1 Introduction
Graph Neural Networks (GNNs) are effective tools for handling graph-structured data and have been widely applied to complex problems such as node classification, graph classification, and edge prediction. They have shown great potential in bioinformatics, chemistry, medicine, cybersecurity, linguistics, and transportation [1]. For example, financial risk control systems can train GNNs on transaction data to identify potential fraudulent activities, such as money laundering [2]. However, graph data are often distributed among multiple owners, making direct access impractical due to privacy concerns, regulations, and competitive dynamics [3].
Federated Learning (FL), as a distributed machine learning method that does not require data sharing, effectively solves the data isolation problem. Training GNNs with FL is known as Federated Graph Neural Networks (FedGNN), which can be categorized into two types: Horizontal Federated Learning (HFGNN), where different participants have the same features but different samples [4, 5]; and Vertical Federated Learning (VFGNN), where different participants have the same samples but different features [6, 7]. This paper focuses on VFGNN, which is more common in heterogeneous data environments such as finance and healthcare. Figure 1 illustrates a FedGNN (VFGNN) collaboration for credit analysis involving a fintech company, an online shopping company, and a social media company. The fintech company holds transaction data and labels, the social media company holds social information, and the online shopping company holds shopping information.
Figure 1: A example of VFGNN backdoor attack
Recently, the security implications of FedGNNs have been brought to the forefront of the community. These popular FedGNN have been proven vulnerable to backdoor attacks [8, 9]. Specifically, backdoor attacks against FedGNN implant hidden backdoors into the victim GNN during federated training, which mislead the model on inputs embedded with trigger patterns while having minimal impact on the clean data during inference. In the scenario shown in Figure 1, the social media company can implant a local backdoor to approve loans for users with poor credit ratings, presenting a significant security threat. There has been active research into investigating backdoor attacks tailored to VFL [10, 11, 12], GNN [13, 14, 15, 16], and HFGNN [9, 8]. However, they are not applicable to VFGNN due to unique trigger pattern restrictions derived from both VFL and GNN. To the best of our knowledge, Backdoor attacks on VFGNN are still unexplored, and the robustness of VFGNN under backdoor attacks is essentially unknown.
To fill this gap, this paper proposes a backdoor attack algorithm for Vertical Federated Graph Neural Network Learning, called BVG. The BVG method generates multi-hop triggers based on the graph structure during VFGNN model training, embedding key features of the target class into the model without accessing or modifying labels. Specifically, our contributions in this paper are summarized as follows.
  • To our best knowledge, we propose the first backdoor attack to reveal the vulnerability of VFGNN.
  • We design multi-hop triggers for the graph structure, which inject triggers into the VFGNN without modifying the graph structure.
  • We design a bi-level optimization trigger generation method, which generates triggers based on the graph structure during VFGNN model training. This process creates a trigger containing key features of the target class and integrates it into the model during VFGNN training.
  • Through extensive experiments on widely used datasets, we demonstrate the effectiveness of our proposed method. The BVG method requires knowledge of only 4 target class nodes to achieve a high success rate of backdoor attacks.
@@ 2 Related Work
2.1 Backdoor attacks on Vertical Federated Learning
@@
Federated Learning is especially vulnerable to backdoor attacks due to its distributed nature [17]. However, the vertically split model of VFL restricts the attacker’s access to the training labels, let alone modify them.
@@
There are two notable directions in attempting to solve this issue. The first is to conduct label inference attacks [18] to get some labels in advance. For example, VILLAIN [19] leverages label inference to pinpoint samples of the target label and then poisons these samples to inject the backdoor. BadVFL [10] and LR-BA [20] assume attackers can acquire labels of a certain number of samples from each category with label inference. Another direction is to make the knowledge of a small number of labeled samples from the target class a prerequisite [21, 11, 12]. These methods try to use as few as possible (e.g., 500 for [21] and 0.1% for [11]) the labeled target-class samples to establish links between backdoor triggers and target label. Normally, in VFL, the backdoor attack can only be conducted in a clean-label manner [22]. Therefore, works like TPGD [23] assuming the label modification capability of the active party are impractical.
2.2 Backdoor attacks on Graph Neural Networks
@@
The unique challenge to graph-oriented backdoor attacks lies in the inherently unstructured and discrete nature of graph data [14]. For structured and continuous data like images, attackers can directly stamp a trigger pattern (e.g., a black square on the bottom right of the image) $s$ onto a benign image $x$ to achieve a poisoned sample $x_{p}=x+s$ [13]. However, the backdoor attacks against GNN involve design triggers within a large spectrum, including topological structures and descriptive (nodes and edges) features [14]. The effectiveness of backdoor attacks largely depends on designing triggers tailored to the specific attributes of the target task. These triggers mainly fall into three categories: malicious subgraph structures, graph structure perturbations, and node attribute manipulation.
@@
Malicious subgraph structures involve adding malicious subgraphs to nodes [24], edge endpoints [16], or specific positions in the graph [25, 14], causing nodes, edges, or graphs to be misclassified. Graph structure perturbation involves injecting special node connections into the training set adversarially, allowing backdoor attacks to be executed by merely modifying the graph’s topology during the attack [26]. Node attribute manipulation involves carefully designing special node features and adding them to the target class nodes in the training set, followed by adaptive optimization of the graph structure to train a GNN model with backdoors [15, 27]. Although these methods are viable in centralized GNN scenarios, they all rely on full access to the training set, which is nearly impossible in VFGNN.
@@
The aforementioned backdoor techniques, whether tailored for GNN or VFL, are not applicable to VFGNN due to unique threat models. Existing methods such as CBA, DBA, and Bkd-FedGNN focus on HFGNN and do not address the needs of VFGNN [9, 8]. VFL backdoor attack methods lack consideration for the data structure of graph data, while GNN backdoor attacks face difficulties in handling vertically partitioned data. To address these issues, we propose the BVG algorithm for VFGNN, which uses dual-layer optimization and multi-hop triggers to achieve high attack success rates with minimal impact on the main task accuracy (MTA).
@@ 3.1 Problem Formulation
In VFGNN, $K$ ($K>=2$) participants collaboratively train a model using their private data. Each participant $k$ has its local graph $G^{k}=(V,E^{k},X^{k})$, which comprises the entire node set $V$ a subset of edges $E^{k}$, and a subset of feature data $X^{k}$ [6]. From a global perspective, all participants collectively own a global graph $G^{gl}=(V,E^{gl},X^{gl})$, where $E^{gl}=E^{1}\cup\ldots\cup E^{K}$, and for any node attribute in $X$, $x_{k}^{gl}=x_{k}^{1}||\ldots||x_{k}^{K}$, where ||||| | is the concatenation operator. One of the $K$ participants is an active party that knows the corresponding labels of the labeled node set $V_{L}\subseteq V$, and all other participants are negative parties that have no access to the label information.
Each participant employs a local GNN model $f_{k}$ parameterized by $\theta_{k}$ to compute the local output $H^{k}_{l}=f_{k}(\mathcal{G}_{l}^{k};\theta_{k})$, where $\mathcal{G}_{l}^{k}$ denotes the computation graph of node $v_{l}$. In addition to training its own local model, the active party also trains a top model $G$ parameterized by $\theta_{top}$, which aggregates the outputs from all local models to minimize the loss function $\mathcal{L}$. Therefore, the VFGNN model training can be formulated as
@@
As an adversary in a backdoor attack, the objective is to establish a mapping association between a trigger and the backdoor target class. This is achieved by injecting a backdoor during the VFGNN training process. Subsequently, upon deployment of the VFGNN model, the adversary can introduce the local trigger to deliberately misclassify the corresponding sample as the target class. The backdoor attack objective in VFGNN is
@@
where $F$ refers to the VFGNN model, which encompasses both the top model and the bottom models of all participating parties. $\delta$ represents the backdoor trigger, a(·) denotes the operation of trigger attachment, and $\tau$ is the backdoor target class.
@@ 3.2 Threat model
Adversary’s objective. In VFGNN multi-classification tasks, the adversary’s objective is to inject a backdoor into the model during the training phase. When the poisoned model is deployed for applications, it will misclassify any local data with the backdoor trigger as the designated target class but maintains classification capacity for all other clean data.
Adversary’s knowledge. In backdoor attacks, the adversary requires only a very few training samples labeled target class. Although this requirement deviates from the original VFGNN setting, it is feasible in practical VFGNN applications because it is possible for an adversary to acquire a small number of target class samples through various means, such as direct purchasing [18, 20]. Our experiment results indicate that acquiring only one to four target class nodes is sufficient for a successful backdoor attack. Apart from these limited target samples, the adversary has no information about the models and data of other parties.
Figure 2: A sketch of our proposed Backdoor Attack for VFGNN.
@@ 3.3 Multi-hop Trigger Generation
Due to the special nature of the VFGNN architecture, all participants are aware of all nodes in the dataset, but possess different attributes of the nodes. Therefore, the adversary cannot generate a subgraph structure as a trigger like common GNN backdoor attacks [24, 28]. Here, the trigger is added to the node attributes. To prepare a trigger, we propose a multi-hop trigger generation method without introducing new nodes. The trigger is generated and trained with the VFGNN model using bi-level optimization. We use the PGD method [29, 30] to generate the trigger. The optimization formula for the trigger is as follows:
@@
where the trigger $\delta$ is trained according to all known target class nodes $\mathcal{V}_{P}$ available to the adversary, $\mathcal{V}_{P}\in\mathcal{V}_{L}$. Essentially, this trigger serves as a universal trigger for the target class $\tau$ [31, 32].
@@
0:  Training dataset $\mathcal{G}$; number of MTG epochs $\mathcal{T}$; target class nodes $\mathcal{V}_{P}$ with target label $\tau$
@@
3.4 Backdoor Trigger and Deployment
After the training is completed, the attacker has successfully injected a backdoor into the model. When the VFGNN model is deployed in practical applications, any local data containing the trigger (added using the method described in (4)) will be misclassified as the specified target class, while other clean data will still be correctly classified. The attacker successfully implements the backdoor attack using a small number of labeled target class samples.
@@ 4.1.1 Local GNN Models
Graph Convolutional Network (GCN) [33]: GCN captures local features by propagating and aggregating information between nodes and their neighbors. Each local GNN model in VFGNN uses a two-layer GCN.
Graph Attention Network (GAT) [34]: GAT uses an attention mechanism to assign adaptive weights to neighbor nodes, enhancing the model’s ability to focus on important nodes.
Simple Graph Convolution (SGC) [35]: SGC simplifies GCN by removing nonlinear activations and reducing layers, making it more efficient while retaining essential graph structural information.
To ensure a fair comparison with previous studies [36], we use a two-layer GNN model for each participant’s local GNN to extract local node embeddings, with the dimension set to 16. The number of hidden units is fixed at 32. For GCN and GAT, the activation function is ReLU. VFGNN is trained using Adam, with a learning rate of 0.01.

### v2

@@ Abstract
Federated Graph Neural Network (FedGNN) integrate federated learning (FL) with graph neural networks (GNNs) to enable privacy-preserving training on distributed graph data. Vertical Federated Graph Neural Network (VFGNN), a key branch of FedGNN, handles scenarios where data features and labels are distributed among participants. Despite the robust privacy-preserving design of VFGNN, we have found that it still faces the risk of backdoor attacks, even in situations where labels are inaccessible. This paper proposes BVG, a novel backdoor attack method that leverages multi-hop triggers and backdoor retention, requiring only four target-class nodes to execute effective attacks. Experimental results demonstrate that BVG achieves nearly 100% attack success rates across three commonly used datasets and three GNN models, with minimal impact on the main task accuracy. We also evaluated various defense methods, and the BVG method maintained high attack effectiveness even under existing defenses. This finding highlights the need for advanced defense mechanisms to counter sophisticated backdoor attacks in practical VFGNN applications.
@@ 1 Introduction
Graph Neural Networks (GNNs), with their powerful capability to process graph-structured data, have demonstrated significant value in cross-domain applications such as bioinformatics, chemical analysis, medical diagnosis, and financial risk management Wu et al. (2020). Specifically, in financial fraud detection scenarios, the use of transaction relationship graphs to identify potential fraudulent activities highlights the unique advantages of GNNs Cheng et al. (2023). However, in real-world settings, graph data is often distributed across different stakeholders, and due to data privacy regulations and the need to protect commercial secrets, traditional centralized training paradigms face severe challenges Liu et al. (2024).
Federated Graph Neural Networks (FedGNN), as a combination of Federated Learning (FL) and GNNs, provide an innovative solution to this challenge He et al. (2021). Among them, Vertical Federated Graph Neural Networks (VFGNN) demonstrate unique value in cross-organizational collaboration scenarios due to their ability to handle heterogeneous distributions of features and labels Mai and Pang (2023). For instance, as shown in Figure 1, in a credit evaluation system, financial institutions hold user transaction records and credit labels, social media platforms possess social graphs, and e-commerce platforms provide consumer behavior features. This collaborative multi-party data setting is a typical application scenario for VFGNN.
Figure 1: An example of VFGNN backdoor attack
Although VFGNN effectively protects privacy in distributed scenarios, it also faces significant security challenges, with backdoor attacks being a major potential threat. Attackers can implant concealed trigger patterns to induce specific inputs to produce predefined malicious outputs while maintaining the model’s normal predictive performance, posing severe risks to the practical application of VFGNN. For example, in the credit evaluation scenario illustrated in Figure 1, malicious participants could use backdoor attacks to enable high-risk users to obtain credit approvals, leading to serious financial security issues. However, research on backdoor attacks targeting VFGNN remains nearly nonexistent. This is primarily due to two major challenges faced by traditional backdoor attacks in VFGNN scenarios: first, the label isolation mechanism among participants restricts traditional attack paths based on label manipulation; second, the complex topology of graph data requires backdoor triggers to not only maintain structural consistency but also satisfy concealment requirements.
To address the above challenges, this paper proposes a novel graph-structured backdoor attack method based on multi-hop adjacency relationships, termed BVG. This method innovatively designs a backdoor injection mechanism tailored for VFGNNs. First, it employs a multi-hop adjacency trigger generation algorithm to covertly propagate target-class features within the graph structure. Second, a backdoor retention strategy is adopted to enhance the attack’s stability. Experimental results demonstrate that BVG requires knowledge of only four target-class nodes to achieve an attack success rate of over 99%, with minimal impact on the performance of the main task. The main contributions of this paper are as follows:
  • Reveal backdoor attack risks in VFGNNs: A systematic analysis of security threats within the VFGNN framework highlights unique risk characteristics distinct from traditional scenarios.
  • Design multi-hop graph structural triggers: A bi-level optimization strategy generates backdoor triggers for graph propagation, enabling flexible and covert backdoor injection without disrupting the graph structure.
  • Develop a backdoor retention strategy: This effectively addresses instability issues in backdoor injection during federated training, improving attack persistence.
  • Validate effectiveness through experiments: Extensive experiments on widely-used benchmark datasets demonstrate that BVG is efficient and stable, maintaining high attack success rates even under various backdoor defense mechanisms.
@@ 2 Related Work
2.1 Backdoor attacks on VFL
@@
Federated Learning is especially vulnerable to backdoor attacks due to its distributed nature Zhang et al. (2024). However, the vertically split model of VFL restricts the attacker’s access to the training labels, let alone modify them.
@@
There are two notable directions in attempting to solve this issue. The first is to conduct label inference attacks Fu et al. (2022) to get some labels in advance. For example, VILLAIN Bai et al. (2023) leverages label inference to pinpoint samples of the target label and then poisons these samples to inject the backdoor. BadVFL Naseri et al. (2024) and LR-BA Gu and Bai (2023) assume attackers can acquire labels of a certain number of samples from each category with label inference. Another direction is to make the knowledge of a small number of labeled samples from the target class a prerequisite Chen et al. (2024, 2023a); He et al. (2023). These methods try to use as few as possible (e.g., 500 for Chen et al. (2024) and 0.1% for Chen et al. (2023a)) the labeled target-class samples to establish links between backdoor triggers and target labels. Normally, in VFL, the backdoor attack can only be conducted in a clean-label manner Zhao et al. (2020). Therefore, works like TPGD Liu et al. (2022) assuming the label modification capability of the active party are impractical.
2.2 Backdoor attacks on GNN
@@
The unique challenge to graph-oriented backdoor attacks lies in the inherently unstructured and discrete nature of graph data Xi et al. (2021). For structured and continuous data like images, attackers can directly stamp a trigger pattern (e.g., a black square on the bottom right of the image) $s$ onto a benign image $x$ to achieve a poisoned sample $x_{p}=x+s$ Li et al. (2022a). However, the backdoor attacks against GNN involve design triggers within a large spectrum, including topological structures and descriptive (nodes and edges) features Xi et al. (2021). The effectiveness of backdoor attacks largely depends on designing triggers tailored to the specific attributes of the target task. These triggers mainly fall into three categories: malicious subgraph structures, graph structure perturbations, and node attribute manipulation.
@@
Malicious subgraph structures involve adding malicious subgraphs to nodes Dai et al. (2023), edge endpoints Zheng et al. (2023), or specific positions in the graph Zhang et al. (2021); Xi et al. (2021), causing nodes, edges, or graphs to be misclassified. Graph structure perturbation involves injecting special node connections into the training set adversarially, allowing backdoor attacks to be executed by merely modifying the graph’s topology during the attack Yang et al. (2022). Node attribute manipulation involves carefully designing special node features and adding them to the target class nodes in the training set, followed by adaptive optimization of the graph structure to train a GNN model with backdoors Xing et al. (2023); Chen et al. (2023b). Although these methods are viable in centralized GNN scenarios, they all rely on full access to the training set, which is nearly impossible in VFGNN.
@@
The aforementioned backdoor techniques, whether tailored for GNN or VFL, are not applicable to VFGNN due to unique threat models. Existing methods such as CBA, DBA, and Bkd-FedGNN focus on HFGNN and do not address the needs of VFGNN Xu et al. (2022); Liu et al. (2023). VFL backdoor attack methods lack consideration for the data structure of graph data, while GNN backdoor attacks face difficulties in handling vertically partitioned data. To address these issues, this paper proposes the BVG algorithm for VFGNN, leveraging multi-hop triggers and a backdoor retention strategy to achieve efficient backdoor attacks.
@@ 3 Proposed Approach
Figure 2: A sketch of our proposed Backdoor Attack for VFGNN.
@@ 3.1 Problem Formulation
In VFGNN, $K$ ($K>=2$) participants collaboratively train a model using their private data. Each participant $k$ has its own local graph $G^{k}=(V,E^{k},X^{k})$, which comprises the entire node set $V$, a subset of edges $E^{k}$, and a subset of feature data $X^{k}$ Chen et al. (2020). From a global perspective, all participants collectively own a global graph $G^{gl}=(V,E^{gl},X^{gl})$, where $E^{gl}=E^{1}\cup\ldots\cup E^{K}$, and for any node attribute in $X$, $x_{i}^{gl}=x_{i}^{1}||\ldots||x_{i}^{K}$, where ||||| | is the concatenation operator. One of the $K$ participants is an active party that knows the corresponding labels of the labeled node set $V_{L}\subseteq V$, and all other participants are negative parties that have no access to the label information.
Each participant employs a local GNN model $f_{k}$ parameterized by $\theta_{k}$ to compute the local output $H^{k}_{i}=f_{k}(\mathcal{G}_{i}^{k};\theta_{k})$, where $\mathcal{G}_{i}^{k}$ denotes the computation graph of node $v_{i}$. In addition to training its own local model, the active party also trains a top model $G$ parameterized by $\theta_{top}$, which aggregates the outputs from all local models to minimize the loss function $\mathcal{L}$. Therefore, the VFGNN model training can be formulated as
@@
As an adversary in a backdoor attack, the main objectives are threefold: first, to ensure that the federated learning task can be completed successfully; second, to establish a mapping between the trigger and the backdoor target class; and third, to ensure that the backdoor injection remains undetected by the active party. These three objectives must be achieved simultaneously during the VFGNN training process. Specifically, the attacker needs to inject the backdoor cleverly, ensuring that the training process is not disrupted and that no suspicious behavior is exposed. After the VFGNN model is deployed, the attacker can introduce a local trigger to deliberately misclassify the corresponding sample as the target class, thereby completing the backdoor attack. The backdoor attack objective in VFGNN is
@@
$+\underbrace{\sum_{v_{i}\in V}||\tilde{H}_{i}-H_{i}||^{2}}_{\text{Undetected Task}}$
where $F$ refers to the VFGNN model, which encompasses both the top model and the bottom models of all participating parties. $\delta$ represents the backdoor trigger, a(·) denotes the operation of trigger attachment, and $\tau$ is the backdoor target class. $\tilde{H}$ denotes the output of the bottom GNN model after the trigger injection, while $H$ refers to the output without trigger injection.
@@ 3.2 Threat model
Adversary’s objective. In VFGNN multi-classification tasks, the adversary’s objective is to inject a backdoor into the model during the training phase without being detected. When the poisoned model is deployed for applications, it will misclassify any local data with the backdoor trigger as the designated target class but maintains classification capacity for all other clean data.
Adversary’s knowledge. In backdoor attacks, the adversary requires only a very few training samples labeled target class. Although this requirement deviates from the original VFGNN setting, it is feasible in practical VFGNN applications because it is possible for an adversary to acquire a small number of target class samples through various means, such as direct purchasing Fu et al. (2022); Gu and Bai (2023). Our experiment results indicate that acquiring only one to four target class nodes is sufficient for a successful backdoor attack. Apart from these limited target samples, the adversary has no information about the models and data of other parties.
@@ 3.3 Multi-hop Trigger Generation
Due to the special nature of the VFGNN architecture, all participants are aware of all nodes in the dataset, but possess different attributes of the nodes. Therefore, the adversary cannot generate a subgraph structure as a trigger like common GNN backdoor attacks Dai et al. (2023); Wang et al. (2024). Here, the trigger is added to the node attributes. To prepare a trigger, we propose a multi-hop trigger generation method without introducing new nodes. The trigger generation and training are conducted synchronously with the VFGNN model. Additionally, to enhance optimization efficiency, we introduce a hyperparameter $\epsilon$ as a trigger threshold. By appropriately setting this threshold, we ensure the stealthiness of the trigger while simplifying the optimization task from a triple optimization to a dual optimization. We use the PGD method Madry et al. (2018); Huang et al. (2021) to generate the trigger. The optimization formula for the trigger is as follows:
@@
where the trigger $\delta$ is trained according to all known target class nodes $\mathcal{V}_{P}$ available to the adversary, $\mathcal{V}_{P}\in\mathcal{V}_{L}$. Essentially, this trigger serves as a universal trigger for the target class $\tau$ Shafahi et al. (2020); Zeng et al. (2022).
@@
0:  Training dataset $\mathcal{G}$; number of MTG epochs $T$; target class nodes $\mathcal{V}_{P}$ with target label $\tau$
@@
3.4 Backdoor Retention
When injecting backdoors, using only a small number of samples may lead to unstable effects. For instance, a backdoor may perform well in one epoch but degrade suddenly in the next. To ensure the stability of backdoors, we propose a method called Backdoor Retention (BR).
The BM method consists of two key steps. First, the attacker evaluates the impact of the backdoor. Since the attacker cannot directly access the output of the VFGNN model during training, they rely on an assumption: if the backdoor is successfully injected, the outputs $H$ of nodes triggered by the injection should exhibit high similarity. Based on this assumption, the backdoor effectiveness $E$ is approximated using the following formula:
$E=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{H_{i}\cdot H_{j}}{\|H_{i}\|_{2}\|H_{j}\|_{2}},$ (6)
where $n$ is the number of nodes involved in the trigger injection, and $H_{i}$ represents the output of node $v_{i}$ in the attacker’s bottom model.
After estimating $E$, the attacker performs the following actions in each epoch: if $E$ exceeds a predefined threshold, the attacker updates the bottom model and trigger; otherwise, the model and trigger from the previous epoch are retained. This approach ensures the stability and effectiveness of the backdoor throughout the training process.
@@ 4.1.1 Local GNN Models
Graph Convolutional Network (GCN) Kipf and Welling (2016): GCN captures local features by propagating and aggregating information between nodes and their neighbors. Each local GNN model in VFGNN uses a two-layer GCN.
Graph Attention Network (GAT) Veličković et al. (2017): GAT uses an attention mechanism to assign adaptive weights to neighbor nodes, enhancing the model’s ability to focus on important nodes.
Simple Graph Convolution (SGC) Wu et al. (2019): SGC simplifies GCN by removing nonlinear activations and reducing layers, making it more efficient while retaining essential graph structural information.
To ensure a fair comparison with previous studies Chen et al. (2022), we use a two-layer GNN model for each participant’s local GNN to extract local node embeddings, with the dimension set to 16. The number of hidden units is fixed at 32. For GCN and GAT, the activation function is ReLU. VFGNN is trained using Adam, with a learning rate of 0.01. We provide a detailed description of the experimental details in the appendix A.